# MusicMagus: Zero-Shot Text-to-Music Editing via Diffusion Models

### v2

@@ Abstract
Recent advances in text-to-music generation models have opened new avenues in musical creativity. However, music generation usually involves iterative refinements, and how to edit the generated music remains a significant challenge. This paper introduces a novel approach to the editing of music generated by such models, enabling the modification of specific attributes, such as genre, mood and instrument, while maintaining other aspects unchanged. Our method transforms text editing to latent space manipulation while adding an extra constraint to enforce consistency. It seamlessly integrates with existing pretrained text-to-music diffusion models without requiring additional training. Experimental results demonstrate superior performance over both zero-shot and certain supervised baselines in style and timbre transfer evaluations. We also show the practical applicability of our approach in real-world music editing scenarios.  [^1] Code and demo: https://bit.ly/musicmagus-demo.  [^2] Work done during Yixiao’s internship at Sony AI.
@@ 1 Introduction
Recent advances in text-to-music generation have unlocked new possibilities in musical creativity Zhang et al. (2020); Lu et al. (2023); Min et al. (2023); Agostinelli et al. (2023); Schneider et al. (2023); Copet et al. (2023); Chen et al. (2023). However, a significant challenge persists in how to edit the generated results as music production usually involves iterative refinements. Building on this momentum, we regard ‘text-to-music editing’ as the process of using text queries to edit music, and we see two major types of operations: inter-stem editing, such as adding or removing instruments (e.g., “add a saxophone” or “remove the drums”), and intra-stem editing, which involves modifications within the same stem, such as adding effects or changing instruments (e.g., “add reverb to this stem” or “transfer the timbre of the specified notes”). In this context, a “stem” refers to an individual track or component within a music piece, such as a specific instrument or vocal part. The primary focus of this paper is on the latter, intra-stem editing.
@@
    Comparative experiments validate that MusicMagus outperforms existing zero-shot methods and some supervised approaches in critical tasks such as style and timbre transformation.
@@ 3 Background
MusicMagus utilizes a pretrained diffusion model Ho et al. (2020) for text-to-music editing, eliminating the need for additional training. Specifically, we use a pretrained AudioLDM 2 model Liu et al. (2023a) as the backbone model, since it is the most accessible text-to-music diffusion model with publicly available weights. AudioLDM 2 employs a variational autoencoder (VAE) Kingma and Welling (2013) to compress a music audio spectrogram into a latent low-dimensional space. It then trains a latent diffusion model (LDM) on this latent space to generate new samples from Gaussian noise conditioned on text inputs. During generation, the LDM takes a condition $y$, generates a latent variable $z_{0}$, and uses the VAE decoder to produce the music spectrogram $x$. This spectrogram can then be converted into a waveform using an external vocoder, such as HiFi-GAN Su et al. (2020).
@@ 4.1 Finding editing direction
In this section, we introduce a strategy to calculate a difference ($\Delta$) vector in the latent space to guide the editing direction.
This method is chosen over direct word swapping as it better preserves semantic coherence and contextual relevance, especially in cases of varying phrase lengths and complex content alterations. We will further explain it in Section 4.2; besides, previous research finds that similar operations can facilitate a more robust edit, especially when the keywords subject to modification are sparsely represented in the training dataset Parmar et al. (2023).
@@
$\epsilon_{\theta}=\epsilon_{\theta}(z_{t},E,t).$ (6)
@@ 5.2 Metrics
Overall Quality (OVL): This metric is used to assess the overall music quality, encompassing aspects like sound clarity and musicality. It primarily evaluates whether the editing process enhances or diminishes the quality of the original music audio. The scoring for this metric ranges from 0 to 100.
Relevance (REL): REL measures the perceived semantic closeness between the edited music and the new text prompt. It is a subjective score, also ranging from 0 to 100.
Structural Consistency (CON): We define a new metric CON to evaluate the consistency of the pitch contour and structural aspects in the subjective test. Similar to the others, its scoring range is from 0 to 100.
@@ 6.1 Real music audio editing
MusicMagus offers capabilities for editing real-world music audio, although it is noted that the performance may not match the editing of synthesized music audio generated from diffusion models.
The process begins with the DDIM inversion to estimate the latent representation $\hat{z_{T}}$ of a given real music audio $x$. This step is crucial to facilitate editing with the diffusion model, as depicted in Figure 4.

### v3

@@ Abstract
Recent advances in text-to-music generation models have opened new avenues in musical creativity. However, music generation usually involves iterative refinements, and how to edit the generated music remains a significant challenge. This paper introduces a novel approach to the editing of music generated by such models, enabling the modification of specific attributes, such as genre, mood and instrument, while maintaining other aspects unchanged. Our method transforms text editing to latent space manipulation while adding an extra constraint to enforce consistency. It seamlessly integrates with existing pretrained text-to-music diffusion models without requiring additional training. Experimental results demonstrate superior performance over both zero-shot and certain supervised baselines in style and timbre transfer evaluations. We also showcase the practical applicability of our approach in real-world music editing scenarios.
@@ 1 Introduction
Recent advances in text-to-music generation have unlocked new possibilities in musical creativity Zhang et al. (2020); Zhao and Xia (2021); Lu et al. (2023); Min et al. (2023); Agostinelli et al. (2023); Schneider et al. (2023); Copet et al. (2023); Chen et al. (2023). However, a significant challenge persists in how to edit the generated results as music production usually involves iterative refinements. Building on this momentum, we regard ‘text-to-music editing’ as the process of using text queries to edit music, and we see two major types of operations: inter-stem editing, such as adding or removing instruments (e.g., “add a saxophone” or “remove the drums”), and intra-stem editing, which involves modifications within the same stem, such as adding effects or changing instruments (e.g., “add reverb to this stem” or “transfer the timbre of the specified notes”). In this context, a “stem” refers to an individual track or component within a music piece, such as a specific instrument or vocal part. The primary focus of this paper is on the latter, intra-stem editing.
@@
    Comparative experiments validate that MusicMagus outperforms existing zero-shot methods and some supervised approaches in critical tasks such as style and timbre transformation.  [^1] Work done during Yixiao’s internship at Sony AI.  [^2] Code and demo: https://bit.ly/musicmagus-demo.
@@ 2.2 Text-to-Music Editing
Concurrently, a recent work Manor and Michaeli (2024) also attempts the similar task of text-based audio editing with diffusion models. We encourage the readers to refer to the work for more details.
@@ 3 Background
MusicMagus utilizes a pretrained diffusion model Ho et al. (2020) for text-to-music editing, eliminating the need for additional training. Specifically, we use a pretrained AudioLDM 2 model Liu et al. (2023a) as the backbone model, since it is the most accessible text-to-music diffusion model with publicly available weights. AudioLDM 2 employs a variational autoencoder (VAE) Kingma and Welling (2013) to compress a music audio spectrogram into a latent low-dimensional space. It then trains a latent diffusion model (LDM) on this latent space to generate new samples from Gaussian noise conditioned on text inputs. During generation, the LDM takes a condition $y$, generates a latent variable $z_{0}$, and uses the VAE decoder to produce the music spectrogram $x$, which can be then converted into a waveform using an external vocoder, such as HiFi-GAN Su et al. (2020).
@@ 4.1 Finding Editing Direction
In this section, we introduce a strategy to calculate a difference ($\Delta$) vector in the latent space to guide the editing direction. This method is chosen over direct word swapping as it better preserves semantic coherence and contextual relevance, especially in cases of varying phrase lengths and complex content alterations. We will further explain it in Section 4.2; besides, previous research finds that similar operations can facilitate a more robust edit, especially when the keywords subject to modification are sparsely represented in the training dataset Parmar et al. (2023).
@@
$\epsilon_{\theta}$ $=\epsilon_{\theta}(z_{t},E,t),$ (6)
@@ 5.2 Metrics
Overall Quality (OVL) This metric is used to assess the overall music quality, encompassing aspects like sound clarity and musicality. It primarily evaluates whether the editing process enhances or diminishes the quality of the original music audio. The scoring for this metric ranges from 0 to 100.
Relevance (REL) REL measures the perceived semantic closeness between the edited music and the new text prompt. It is a subjective score, also ranging from 0 to 100.
Structural Consistency (CON) We define a new metric CON to evaluate the consistency of the pitch contour and structural aspects in the subjective test. Similar to the others, its scoring range is from 0 to 100.
@@ 6.1 Real Music Audio Editing
MusicMagus offers capabilities for editing real-world music audio, although it is noted that the performance may not match the editing of synthesized music audio generated from diffusion models. We begin with the DDIM inversion to estimate the latent representation $\hat{z_{T}}$ of a given real music audio $x$. This step is crucial to facilitate editing with the diffusion model, as depicted in Figure 4.